R&D (FilterCavity)
NaokiAritomi - 14:41, Monday 31 January 2022 (2812)
Two magnets of end mirror fell down

[Takahashi, Aritomi]

Since there is a new 6.35Hz peak in END oplev spectra as reported in elog2808, we opened the end chamber and checked the suspension. We found that two magnets of end mirror fell down! This is the reason why the Z correction does not work recently. Takahashi-san will glue the magnets next Monday.

Comments related to this report
RyutaroTakahashi - 22:46, Monday 31 January 2022 (2815)

Picture of the mirror without magnets.

Images attached to this comment
2815_20220131144514_em2magnet.jpg
RyutaroTakahashi - 22:36, Monday 07 February 2022 (2833)

One magnet (upper) was glued with the jig and released 3 hours later. The other magnet (left) is under gluing. It will be released on the 14th.

Images attached to this comment
2833_20220207143425_magnetonem21.png 2833_20220207143448_magnetonem22.png 2833_20220207143526_magnetonem23.png
RyutaroTakahashi - 11:36, Monday 14 February 2022 (2837)

The other magnet (left) was released. The TM was released and the chamber was closed.

Images attached to this comment
2837_20220214033618_magnetonem24.jpg